A thick walled tube of stainless steel with 20 mm inner diameter and

40 mm outer diameter is covered with a 30

mm layer of asbestos insulation (k = 0.2 W/

m°C). If the inside wall temperature of the pipe

is maintained at 600°C and the outside

insulation at 1000°C, calculate the heat loss

per metre of length.


1
Expert's answer
2021-09-06T06:56:51-0400

Solution;

Given;

"r_1=\frac{20}{2}=10mm=0.01m"

"r_2=\frac{40}{2}=20mm=0.02m"

"r_3=20+30=50mm=0.05m"

"t_1=600\u00b0c"

"t_3=1000\u00b0c"

"k_B=0.2W\/m\u00b0c"

Heat transfer per meter of length is;

From;

"Q=\frac{2πL(t_1-t_3)}{\frac{ln(\frac{r_2}{r_1})}{k_A}+\frac{ln(\frac{r_3}{r_2}}{k_B}}"

Since the thermal conductivity of satinless steel is not given, therefore, neglecting the resistance offered by stainless steel to heat transfer across the tube, we have;

"\frac{Q}{L}=\frac{2π(t_1-t_3)}{\frac{ln(\frac{r_3}{r_2})}{k_B}}"

"\frac{Q}{L}=\frac{2π(600-1000)}{\frac {ln(\frac{0.05}{0.02})}{0.2}}"

"\frac{Q}{L}=-548.57W/m"

Note:

Negative sign indicates that the heat transfer takes place radially inward.